# Physical Effects of Gravitational Waves at Second Order

### v1

@@ Abstract
Primordial fluctuations are the seeds of Cosmic Microwave Background (CMB) anisotropies and large-scale structures in the Universe [27]. The same fluctuations induce GWs at second order, which predict a scale-invariant spectrum $\Omega_{\mathrm{GW}}\sim 10^{-24}$ on large scales [24]. It could become significant on smaller scales, usually related to the primordial black hole scenario [28, 29, 30, 31, 32, 33, 34, 35, 15, 36, 37] and serve as a probe of the Universe’s unexplored expansion history [38, 39, 40, 13, 41]. Yet, there are lingering gauge ambiguities in theoretical predictions for induced GWs even today [20, 42].
@@
where we introduced the measured redshift $1+\tilde{z}\equiv{a_{\rm r}(\tau_{\rm r})}/{a_{\rm e}(\tau_{\rm e}})$ [66], with $\tau$ being the non-linear proper time of emitter and observer. This results coincide with the conventional definition, that is $1+z=(p_{\mu}k_{\gamma}^{\mu})_{\rm r}/(p_{\mu}k_{\gamma}^{\mu})_{\rm e}$, and contains the standard and integrated Sachs-Wolfe, the Doppler and the GW effects [67].
Second-order results.— From Eqs. (3a) and (Physical Effects of Gravitational Waves at Second Order), we see that the GW effect comes from the term proportional to $n^{i}n^{j}$. As general second-order calculations are quite involved (see, e.g., Refs. [68, 69, 70, 71]), we focus solely on terms containing $n^{i}n^{j}$ from now on. These terms lead to the characteristic “quadrupole” nature of GWs.
@@
At this point, let us discuss implications of our results on the time-independent GW strain induced by the constant gravitational potential $\Phi^{(1)}$ in matter dominated universes [74, 24, 75, 42, 50, 76, 77]. This constant solution, if generated in early matter dominated phases, has been shown to be very sensitive to the transition to the radiation dominated phase [78, 79]. Looking at Eq. (11), it is clear that this constant mode does not affect the redshift. Therefore, GW detectors using time series (like PTAs and Doppler tracking) will not see such a time-independent strain. The same applies to CMB temperature fluctuations and B-modes [80, 81, 82, 83] as they also depend on the redshift. Furthermore, it does not affect the GW energy density if we consider a definition in which $\rho_{\rm GWs}\propto\langle h_{ij}^{\prime}h_{ij}^{\prime}\rangle$ as proposed by Refs. [51, 77].
@@
$\tau^{(2)}_{\rm rf}\supset a_{\rm f}\,$ $\frac{1}{2}n^{i}n^{j}\int_{0}^{L}{\rm d}\lambda\,h^{N(2)}_{ij}(x^{i})$
$=a_{\rm f}\,\frac{1}{2}n^{i}n^{j}\int_{0}^{L}n_{k}{\rm d}x^{k}\,h^{N(2)}_{ij}(x^{i})\,$, (12)
where we used that since $h^{N(2)}_{ij}(x^{i})$ does not depend on time $\int d\lambda=\int n_{k}{\rm d}x^{k}$. The resulting effect is that our sphere of detectors is now displaced by $\Delta L_{\rm i}(x^{i})=\frac{1}{2}n^{i}n^{j}h^{N(2)}_{ij}(x^{i})$; see Eq. (6) and the surrounding discussion. If we calibrate the detector after generation of the constant strain, we eliminate the memory effect. However, if our geodesic clock was calibrated before strain generation, we should be able to see the strain build-up as in the GW memory effect. Moreover, since the constant strain produces a non-trivial Riemann tensor, it may have non-local effects. Thus, despite Ref. [50] showed that the constant mode can be locally gauged away, it is conceivable that it could be observed by extended measurements.
Conclusion.—In this work, we computed the time delay caused by passing GWs at second order in cosmological perturbation theory as measured by geodesic observers exchanging EM signals, without gauge fixing. We showed that the time delay measured by a GW detector is gauge invariant. Then, we identified the effect of passing GWs as that with a quadrupolar dependence on the direction of propagation of the EM signal. We found that, after removing ambiguities associated to the motion of the observers, the resulting physical GW effect to the time delay at second order coincides with the TT components of the spatial metric evaluated in the Newtonian gauge. Therefore, the gauge issue of the induced GW strain is solved.
@@
We have only focused on scalar squared contributions to the second-order transverse-traceless components, since they are the relevant quantities for the (scalar-scalar) induced GWs. But in general, there are also scalar-tensor contributions. In fact, it has been found that scalar-tensor induced GW suffer from divergences in some cases [86, 87, 88, 89]. If it is a gauge issue, it would be clear in a generalization of the time delay calculation to include scalar-tensor terms. It would also be interesting to study the geodesic deviation equation in terms of local tetrads and confirm that $h_{ij}^{N(2)}$ (9) is the quantity describing the effect of GWs at second order.
We conclude by clarifying that, although we demonstrated that the GW strain measured by a GW detector coincides with one in the Newton gauge, our result does not imply that the energy density associated with GWs at second order coincides with the one computed in the Newton gauge. Namely, there is an additional gauge ambiguity in the definition of the (pseudo) energy momentum tensor of GWs [90] at second order in cosmological perturbation theory (see, e.g., Ref. [51] for a study in a fixed time-slice). This is, however, a different problem. It concerns the backreaction of second-order perturbations on the background spacetime, which is important for estimating the expansion rate of the Universe, but is unrelated to the strain measured by a GW detector. We leave this issue for future work.

### v2

@@ Abstract
Primordial fluctuations are the seeds of Cosmic Microwave Background (CMB) anisotropies and large-scale structures in the Universe [27]. The same fluctuations induce GWs at second order, which predict a scale-invariant spectrum $\Omega_{\mathrm{GW}}\sim 10^{-24}$ on large scales [24]. It could become significant on smaller scales, usually related to the primordial black hole scenario [28, 29, 30, 31, 32, 33, 34, 35, 15, 36, 37, 38] and serve as a probe of the Universe’s unexplored expansion history [39, 40, 41, 13, 42]. Yet, there are lingering gauge ambiguities in theoretical predictions for induced GWs even today [20, 43].
@@
where we introduced the measured redshift $1+\tilde{z}\equiv{a_{\rm r}(\tau_{\rm r})}/{a_{\rm e}(\tau_{\rm e}})$ [67], with $\tau$ being the non-linear proper time of emitter and observer. This results coincide with the conventional definition, that is $1+z=(p_{\mu}k_{\gamma}^{\mu})_{\rm r}/(p_{\mu}k_{\gamma}^{\mu})_{\rm e}$, and contains the standard and integrated Sachs-Wolfe, the Doppler and the GW effects [68, 69].
Second-order results.— From Eqs. (3a) and (Physical Effects of Gravitational Waves at Second Order), we see that the GW effect comes from the term proportional to $n^{i}n^{j}$. As general second-order calculations are quite involved (see, e.g., Refs. [70, 71, 72, 73, 74, 75, 76, 77]), we focus solely on terms containing $n^{i}n^{j}$ from now on. These terms lead to the characteristic “quadrupole” nature of GWs.
@@
At this point, let us discuss implications of our results on the time-independent GW strain induced by the constant gravitational potential $\Phi^{(1)}$ in matter dominated universes [80, 24, 81, 43, 51, 82, 83]. This constant solution, if generated in early matter dominated phases, has been shown to be very sensitive to the transition to the radiation dominated phase [84, 85]. Looking at Eq. (11), it is clear that this constant mode does not affect the redshift. Therefore, GW detectors using time series (like PTAs and Doppler tracking) will not see such a time-independent strain. The same applies to CMB temperature fluctuations and B-modes [86, 87, 88, 89] as they also depend on the redshift. Furthermore, it does not affect the GW energy density if we consider a definition in which $\rho_{\rm GWs}\propto\langle h_{ij}^{\prime}h_{ij}^{\prime}\rangle$ as proposed by Refs. [90, 52, 83].
@@
$\tau^{(2)}_{\rm rf}\supset a_{\rm rf}\,$ $\frac{1}{2}n^{i}n^{j}\int_{0}^{L}{\rm d}\lambda\,h^{N(2)}_{ij}(x^{i})$
$=a_{\rm rf}\,\frac{1}{2}n^{i}n^{j}\int_{0}^{L}n_{k}{\rm d}x^{k}\,h^{N(2)}_{ij}(x^{i})\,$, (12)
where we used that since $h^{N(2)}_{ij}(x^{i})$ does not depend on time $\int d\lambda=\int n_{k}{\rm d}x^{k}$. The resulting effect is that our sphere of detectors is now displaced by $\Delta L_{\rm i}(x^{i})=\frac{1}{2}\int_{0}^{L}n_{k}{\rm d}x^{k}n^{i}n^{j}h^{N(2)}_{ij}(x^{i})$; see Eq. (6) and the surrounding discussion. If we calibrate the detector after generation of the constant strain, we eliminate the memory effect. However, if our geodesic clock was calibrated before strain generation, we should be able to see the strain build-up as in the GW memory effect. Moreover, since the constant strain produces a non-trivial Riemann tensor, it may have non-local effects. Thus, despite Ref. [51] showed that the constant mode can be locally gauged away, it is conceivable that it could be observed by extended measurements.
Conclusion.—In this work, we computed the time delay caused by passing GWs at second order in cosmological perturbation theory as measured by geodesic observers exchanging EM signals, without gauge fixing. We showed that the time delay measured by a GW detector is gauge invariant. Then, we identified the effect of passing GWs as that with a quadrupolar dependence on the direction of propagation of the EM signal. We found that, after removing ambiguities associated with the motion of the observers, the resulting physical GW effect on the time delay at second order coincides with the TT components of the spatial metric evaluated in the Newton gauge. Therefore, the gauge issue of the induced GW strain is solved.
Our result also applies to a more general setting, as we have not made any assumptions about the GW source. Namely, we have demonstrated that GW strain predictions from any cosmic source, performed in the Newton gauge, are free from gauge artifacts and directly related to observables.
@@
We have only focused on scalar squared contributions to the second-order transverse-traceless components, since they are the relevant quantities for the (scalar-scalar) induced GWs. But in general, there are also scalar-tensor contributions. In fact, it has been found that scalar-tensor induced GW suffer from divergences in some cases [93, 94, 95, 96, 97]. If it is a gauge issue, it would be clear in a generalization of the time delay calculation to include scalar-tensor terms. It would also be interesting to study the geodesic deviation equation in terms of local tetrads and confirm that $h_{ij}^{N(2)}$ (9) is the quantity describing the effect of GWs at second order.
We conclude by clarifying that, although we demonstrated that the GW strain measured by a GW detector coincides with one in the Newton gauge, our result does not imply that the energy density associated with GWs at second order coincides with the one computed in the Newton gauge. Namely, there is an additional gauge ambiguity in the definition of the (pseudo) energy momentum tensor of GWs [98] at second order in cosmological perturbation theory (see, e.g., Ref. [99, 52] for a study in a fixed time-slice). This is, however, a different problem. It concerns the backreaction of second-order perturbations on the background spacetime, which is important for estimating the expansion rate of the Universe, but is unrelated to the strain measured by a GW detector. We leave this issue for future work.